For the holidays, Marty gave a watch worth $17,300 to Emily and jewelry worth $50,750 to Natalie.

Required: Has Marty made any taxable gifts this year and, if so, in what amounts?

Assume Marty is married to Wendy and they live in a community property state. What are the amounts, if any, of their taxable gifts? (For all requirements, leave no answer blank. Enter zero if applicable.)

a. Martys taxable gift to Emily Martys taxable gift to Natalie Martys taxable gift to Emily Martys taxable gift to Natali
Answer #1

Solution :-

As per the Tax on Gifts in the United States Gifts upto annual exclusion Incurs no tax

The exclusion is different Each year

2002-2005 = $11000

2006 - 2008 = $12000

2009 - 2012 = $13000

2013 - 2015 = $14000

Assume it is 2015 as nothing is given

(A)

After Exclusion of $14000

Marty's taxable gift to Emily = $17300 - $14000 = $2300

Marty's taxable gift to Natalie = $50750 - $14000 = $36750

(B)

In case of Married Couple

Gift Amount distributed Half Half =  ($17500 / 2 ) = $8650

Jwellery Gift half half = $50750 / 2 = $25375

Marty's taxable gift to Emily =( $17500 / 2 ) = $8650 (exclusion upto $14000) = Nil

Marty's taxable gift to Natalie = $25375 - $14000 = $11375

Wendy's taxable gift to Emily =( $17500 / 2 ) = $8650 (exclusion upto $14000) = Nil

Wendy's taxable gift to Natalie = $25375 - $14000 = $11375